# A kernel conditional two-sample test

### v1

@@ D.3.2 Tensor factorization of Hilbert spaces and coefficients transformations
We are now ready to prove Theorem D.14. It does not directly follow from Lemma D.15, but rather its proof is an adaptation of that of this lemma.
Proof of Theorem D.14.
Take $D=\dim({\tilde{\mathcal{H}}})\in{\mathbb{N}_{>0}}\cup\{\infty\}$, $(b_{m})_{m\in\left[D\right]}$@iaciONB ONB of ${\tilde{\mathcal{H}}}$, and a family $(\lambda_{m})_{m\in\left[D\right]}\subset\mathbb{R}_{\geq 0}$ such that $\tilde{V}=\sum_{m=1}^{D}\lambda_{m}b_{m}\otimes b_{m}$, which all exist since $\tilde{V}$ is diagonal. The inverse of $\tilde{V}$ exists and is given by
@@
and observe that $W_{M}$ is of rank $M$ and hence trivially of trace class, and it is self-adjoint and positive semidefinite. Although Lemma D.16 does not directly apply by setting $\tilde{V}^{-1}$ to $W_{M}$ since $W_{M}$ may not be invertible, we see from the proof that the only point where this invertibility is used is in (30). Therefore, the proof carries over until this point, showing that for all $v\in\mathcal{V}$
~$\begin{split}\int_{{\tilde{\mathcal{H}}}}&G_{N}(\tilde{h}\otimes v)\mathrm{d}\mathcal{N}(\tilde{h}\mid 0,W_{M})\\ &=\left[\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\lVert v\rVert^{2}_{\mathcal{V}}\tilde{M}_{N}W_{M}\tilde{M}_{N}^{\star}\right)\right]^{-1/2}\exp\left[\frac{1}{2}\left\lVert\left(\mathrm{id}_{{\tilde{\mathcal{H}}}}+\lVert v\rVert^{2}_{\mathcal{V}}W_{M}^{1/2}\tilde{V}_{N}W_{M}^{1/2}\right)^{-1/2}W_{M}^{1/2}\tilde{S}_{N}(v)\right\rVert_{{\tilde{\mathcal{H}}}}^{2}\right],\end{split}$
@@
$|\langle u,W_{M}u^{\prime}\rangle_{\tilde{\mathcal{H}}}-\langle u,\tilde{V}^{-1}u^{\prime}\rangle_{\tilde{\mathcal{H}}}|$ $=\left|\left\langle u,\sum_{m=1}^{M}\lambda_{m}^{-1}\langle u^{\prime},b_{m}\rangle_{\tilde{\mathcal{H}}}b_{m}\right\rangle-\left\langle u,\sum_{m=1}^{D}\lambda_{m}^{-1}\langle u^{\prime},b_{m}\rangle_{\tilde{\mathcal{H}}}b_{m}\right\rangle\right|$
@@
$\left\lVert\left(\mathrm{id}_{{\tilde{\mathcal{H}}}}+W_{M}^{1/2}\tilde{V}_{N}W_{M}^{1/2}\right)^{-1/2}W_{M}^{1/2}\tilde{S}_{N}(v)\right\rVert_{k}^{2}\to\left\lVert\left(\mathrm{id}_{{\tilde{\mathcal{H}}}}+\tilde{V}^{-1/2}\tilde{V}_{N}\tilde{V}^{-1/2}\right)^{-1/2}\tilde{V}^{-1/2}\tilde{S}_{N}(v)\right\rVert_{k}^{2},$
@@
By the exact same computations as in the proof of Lemma D.18 (which do not rely on whether or not $\tilde{V}^{-1}$ is trace-class), it follows that
$\bar{G}_{N}:=\int_{\mathcal{V}}\bar{G}_{N}(v)\mathrm{d}\mu_{Q}(v)\geq\left[\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}\tilde{V}^{-1}\tilde{M}_{N}^{\star}\right)\right]^{-1/2}\exp\left[\frac{1}{2\mathrm{Tr}(Q)}\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert^{2}_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}\right].$
Furthermore, by construction $\bar{G}_{N}$ is nonnegative. Finally, we have a.s. from Lemma D.19 that
@@
This ensures that $(\bar{G}_{N})_{N}$ is still a supermartingale. Since $\bar{G}_{0}=1$, we also have $\mathbb{E}[\bar{G}_{N}]\leq 1$ for all $N\in{\mathbb{N}_{>0}}$. We can thus leverage Ville’s inequality, and conclude similarly as in the proof of Lemma D.15. ∎
Before concluding this section, we provide an important corollary of Theorem D.14, which provides an upper bound on $\lVert S_{N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes\mathrm{id}_{\mathcal{V}}}$ rather than on $\lVert S_{N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}$ at the price of slightly stronger assumptions on the noise.
D.3.4 Corollary: bound for the noise term of vector-valued least-squares
We now specialize our results to the situation that appears in vector-valued least-squares. The following bound allows to derive a time-uniform high-probability bound for vector-valued least-squares when combined with Lemma D.2, cf. Theorem D.23 below.
Corollary D.20.
Under the setup of Section D.1, assume that Assumptions D.5 and D.13 hold. Let $V\in\mathcal{L}(\mathcal{G})$ be self-adjoint, positive-definite, diagonal, and with bounded inverse. Assume that there exists $\rho\in\mathbb{R}_{>0}$ and $R_{\mathcal{V}}\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and trace-class, such that $\eta$ is $R$-subgaussian conditionally on $\mathcal{F}$, where $R:=\rho^{2}\iota_{\mathcal{G}}(\mathrm{id}_{\tilde{\mathcal{G}}}\otimes R_{\mathcal{V}})\iota_{\mathcal{G}}^{-1}$. Then, for all $\delta\in(0,1)$,
@@ D.3.4 Corollary: bound for the noise term of vector-valued least-squares
Remark D.21.
The subgaussianity assumption of Corollary D.20 is in general a strengthening of the assumption that $\eta$ is $\bar{\rho}$-subgaussian for $\bar{\rho}\in\mathbb{R}_{>0}$, but also a weakening of the one that $\eta$ is $\bar{R}$-subgaussian for $\bar{R}\in\mathcal{L}_{\mathrm{b}}(\mathcal{G})$ self-adjoint, positive semi-definite, and of trace class. Indeed, the operator $R$ given in the corollary is not trace-class when $\tilde{\mathcal{G}}$ is infinite-dimensional. Furthermore, for any such operator $\bar{R}$, there exist $\rho\in\mathbb{R}_{>0}$ and $Q\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and of trace class such that $\bar{R}\preceq R:=\rho^{2}\iota_{\mathcal{G}}(\mathrm{id}_{\tilde{\mathcal{G}}}\otimes Q)\iota_{\mathcal{G}}^{-1}$. Indeed, define the trace-class operator $\bar{R}^{\prime}=\iota_{\mathcal{G}}^{-1}\bar{R}\iota_{\mathcal{G}}\in\mathcal{L}_{\mathrm{b}}({\tilde{\mathcal{H}}}\otimes\mathcal{V})$, and take $\rho=(\lVert\mathrm{Tr}_{\mathcal{V}}(\bar{R}^{\prime})\rVert_{\mathcal{L}_{\mathrm{b}}({\tilde{\mathcal{H}}})})^{1/2}$, and $Q=\mathrm{Tr}_{{\tilde{\mathcal{H}}}}(\bar{R}^{\prime})$, where $\mathrm{Tr}_{W}(A)$ is the partial trace over $W$ of a trace-class operator $A$ defined on a tensor-product space involving $W$. We forego proving this construction for brevity.
The proof of Corollary D.20 relies on the following lemma.
Lemma D.22.
@@
Classical results on scalar subgaussian variables then guarantee that $\mathrm{Var}[X]\leq R_{e}^{2}$. Yet, $e\in\mathrm{ran}(R)^{\perp}=\ker(R^{\star})=\ker(R)$, since $R$ is self-adjoint, and thus $Re=0$. As a result, $R_{e}=\lVert e\rVert_{R}=(\langle Re,e\rangle)^{1/2}=0$, and thus $\mathrm{Var}[X]=0$ and $X$ is constant a.s., and this constant is 00 since $X$ is 00-mean. We deduce by separability of $\mathcal{G}$ and continuity of the scalar product that $\mathbb{P}[\exists e\in\mathrm{ran}(R)^{\perp},~{}\langle\eta,e\rangle\neq 0]=0$, which shows that it holds with full probability that $\eta\in(\mathrm{ran}(R)^{\perp})^{\perp}=\mathrm{cl}(\mathrm{ran}(R))$, concluding the proof. ∎
Proof of Corollary D.20.
Under the current assumptions, the scaled noise $\bar{\eta}_{n}=(R^{\sharp})^{1/2}\eta_{n}$ is $1$-subgaussian conditionally on $\mathcal{F}$, where $R^{\sharp}$ is the Moore-Penrose pseudo-inverse of $R$. Indeed, it is (conditionally) 00-mean and for any $g\in\mathcal{G}$,
@@
The second equality comes from the fact that $R^{1/2}$ and $R^{\sharp}$ commute; the proof is a technical exercise that follows from the spectral theorem applied to $R_{\mathcal{V}}$ and the spectral definition of $R^{\sharp}$. We also leveraged in the last inequality the fact that $RR^{\sharp}$ is an orthogonal projection, and thus $\lVert g\rVert_{RR^{{\sharp}}}\leq\lVert g\rVert_{\mathcal{G}}$; see for instance Proposition 2.3 in Engl et al. [45]. Theorem D.14 applied with $Q=R_{\mathcal{V}}$ then guarantees that for all $\delta\in(0,1)$, it holds with probability at least $1-\delta$ that
$\forall N\in{\mathbb{N}_{>0}},\lVert\iota_{\mathcal{H}}^{-1}M_{N}^{\star}(R^{\sharp}_{N})^{1/2}\eta_{:N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes R_{\mathcal{V}}}\leq\sqrt{2\mathrm{Tr}(R_{\mathcal{V}})\ln\left(\frac{1}{\delta}\left[\det(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}^{\star}\tilde{V}^{-1}\tilde{M}_{N})\right]^{1/2}\right)}.$
@@
[…]
@@
where we leveraged in the third equality the fact that $(R_{\mathcal{V}}^{\sharp})^{1/2}$ and $R_{\mathcal{V}}$ commute, and that $(\tilde{V}+\tilde{V}_{N})^{-1}\otimes\mathrm{id}_{\mathcal{V}}=\iota_{\mathcal{H}}^{-1}(V+V_{N})^{-1}\iota_{\mathcal{H}}$ in the fourth one. Next, $RR^{\sharp}$ is the orthogonal projection on $\mathrm{cl}(\mathrm{ran}(R))$; see for instance Proposition 2.3 in [45]. Yet, $\eta_{n}$ is a.s. in $\mathrm{cl}(\mathrm{ran}(R))$ by Lemma D.22. It follows that $RR^{\sharp}\eta_{n}=\eta_{n}$, a.s., for all $n\in{\mathbb{N}_{>0}}$. As a result, it holds a.s. that
@@
$=\rho^{-2}\lVert M_{N}^{\star}\eta_{:N}\rVert_{(V+V_{N})^{-1}}.$

### v2

@@ E.3.2 Tensor factorization of Hilbert spaces and coefficients transformations
We are now ready to prove Theorem E.14. It does not directly follow from Lemma E.15, but rather its proof is an adaptation of that of this lemma.
Proof of Theorem E.14.
Take $D=\dim({\tilde{\mathcal{H}}})\in{\mathbb{N}_{>0}}\cup\{\infty\}$, $(b_{m})_{m\in\left[D\right]}$ \@iaciONB ONB of $\tilde{\mathcal{H}}$}, and a family $(\lambda_{m})_{m\in\left[D\right]}\subset\mathbb{R}_{\geq 0}$ such that $\tilde{V}=\sum_{m=1}^{D}\lambda_{m}b_{m}\otimes b_{m}$, which all exist since $\tilde{V}$ is diagonal. The inverse of $\tilde{V}$ exists and is given by
@@
and observe that $W_{M}$ is of rank $M$ and hence trivially of trace class, and it is self-adjoint and positive semidefinite. Although Lemma E.16 does not directly apply by setting $\tilde{V}^{-1}$ to $W_{M}$ since $W_{M}$ may not be invertible, we see from the proof that the only point where this invertibility is used is in (30). Therefore, the proof carries over until this point, showing that for all $v\in\mathcal{V}$
$\begin{split}\int_{{\tilde{\mathcal{H}}}}&G_{N}(\tilde{h}\otimes v)\mathrm{d}\mathcal{N}(\tilde{h}\mid 0,W_{M})\\ &=\left[\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\lVert v\rVert^{2}_{\mathcal{V}}\tilde{M}_{N}W_{M}\tilde{M}_{N}^{\star}\right)\right]^{-1/2}\\ &\quad\times\exp\left[\frac{1}{2}\left\lVert\left(\mathrm{id}_{{\tilde{\mathcal{H}}}}+\lVert v\rVert^{2}_{\mathcal{V}}W_{M}^{1/2}\tilde{V}_{N}W_{M}^{1/2}\right)^{-1/2}W_{M}^{1/2}\tilde{S}_{N}(v)\right\rVert_{{\tilde{\mathcal{H}}}}^{2}\right].\end{split}$
@@
[…]
@@
$\left\lVert\left(\mathrm{id}_{{\tilde{\mathcal{H}}}}+W_{M}^{1/2}\tilde{V}_{N}W_{M}^{1/2}\right)^{-1/2}W_{M}^{1/2}\tilde{S}_{N}(v)\right\rVert_{k}^{2}$
$\quad\to\left\lVert\left(\mathrm{id}_{{\tilde{\mathcal{H}}}}+\tilde{V}^{-1/2}\tilde{V}_{N}\tilde{V}^{-1/2}\right)^{-1/2}\tilde{V}^{-1/2}\tilde{S}_{N}(v)\right\rVert_{k}^{2},$
@@
By the exact same computations as in the proof of Lemma E.18 (which do not rely on whether or not $\tilde{V}^{-1}$ is trace-class), it follows that
$\bar{G}_{N}$ $:=\int_{\mathcal{V}}\bar{G}_{N}(v)\mathrm{d}\mu_{Q}(v)$
$\geq\left[\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}\tilde{V}^{-1}\tilde{M}_{N}^{\star}\right)\right]^{-1/2}\exp\left[\frac{1}{2\mathrm{Tr}(Q)}\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert^{2}_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}\right].$
Furthermore, by construction $\bar{G}_{N}$ is nonnegative. Finally, we have a.s. from Lemma E.19 that
@@
This ensures that $(\bar{G}_{N})_{N}$ is still a supermartingale. Since $\bar{G}_{0}=1$, we also have $\mathbb{E}[\bar{G}_{N}]\leq 1$ for all $N\in{\mathbb{N}_{>0}}$. We can thus leverage Ville’s inequality, and conclude similarly as in the proof of Lemma E.15. ∎
Before concluding this section, we provide an important corollary of Theorem E.14, which provides an upper bound on $\lVert S_{N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes\mathrm{id}_{\mathcal{V}}}$ rather than on $\lVert S_{N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}$ at the price of slightly stronger assumptions on the noise.
E.3.4 Corollary: bound for the noise term of vector-valued least-squares
We now specialize our results to the situation that appears in vector-valued least-squares. The following bound allows to derive a time-uniform high-probability bound for vector-valued least-squares when combined with Lemma E.2, cf. Theorem E.23 below.
Corollary E.20.
Under the setup of Section E.1, assume that Assumptions E.5 and E.13 hold. Let $V\in\mathcal{L}(\mathcal{G})$ be self-adjoint, positive-definite, diagonal, and with bounded inverse. Assume that there exists $\rho\in\mathbb{R}_{>0}$ and $R_{\mathcal{V}}\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and trace-class, such that $\eta$ is $R$-subgaussian conditionally on $\mathcal{F}$, where $R:=\rho^{2}\iota_{\mathcal{G}}(\mathrm{id}_{\tilde{\mathcal{G}}}\otimes R_{\mathcal{V}})\iota_{\mathcal{G}}^{-1}$. Then, for all $\delta\in(0,1)$,
@@ E.3.4 Corollary: bound for the noise term of vector-valued least-squares
Remark E.21.
The subgaussianity assumption of Corollary E.20 is in general a strengthening of the assumption that $\eta$ is $\bar{\rho}$-subgaussian for $\bar{\rho}\in\mathbb{R}_{>0}$, but also a weakening of the one that $\eta$ is $\bar{R}$-subgaussian for $\bar{R}\in\mathcal{L}_{\mathrm{b}}(\mathcal{G})$ self-adjoint, positive semi-definite, and of trace class. Indeed, the operator $R$ given in the corollary is not trace-class when $\tilde{\mathcal{G}}$ is infinite-dimensional. Furthermore, for any such operator $\bar{R}$, there exist $\rho\in\mathbb{R}_{>0}$ and $Q\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and of trace class such that $\bar{R}\preceq R:=\rho^{2}\iota_{\mathcal{G}}(\mathrm{id}_{\tilde{\mathcal{G}}}\otimes Q)\iota_{\mathcal{G}}^{-1}$. Indeed, define the trace-class operator $\bar{R}^{\prime}=\iota_{\mathcal{G}}^{-1}\bar{R}\iota_{\mathcal{G}}\in\mathcal{L}_{\mathrm{b}}({\tilde{\mathcal{H}}}\otimes\mathcal{V})$, and take $\rho=(\lVert\mathrm{Tr}_{\mathcal{V}}(\bar{R}^{\prime})\rVert_{\mathcal{L}_{\mathrm{b}}({\tilde{\mathcal{H}}})})^{1/2}$, and $Q=\mathrm{Tr}_{{\tilde{\mathcal{H}}}}(\bar{R}^{\prime})$, where $\mathrm{Tr}_{W}(A)$ is the partial trace over $W$ of a trace-class operator $A$ defined on a tensor-product space involving $W$. We forego proving this construction for brevity.
The proof of Corollary E.20 relies on the following lemma.
Lemma E.22.
@@
Classical results on scalar subgaussian variables then guarantee that $\mathrm{Var}[X]\leq R_{e}^{2}$. Yet, $e\in\mathrm{ran}(R)^{\perp}=\ker(R^{\star})=\ker(R)$, since $R$ is self-adjoint, and thus $Re=0$. As a result, $R_{e}=\lVert e\rVert_{R}=(\langle Re,e\rangle)^{1/2}=0$, and thus $\mathrm{Var}[X]=0$ and $X$ is constant a.s., and this constant is 0 since $X$ is 0-mean. We deduce by separability of $\mathcal{G}$ and continuity of the scalar product that $\mathbb{P}[\exists e\in\mathrm{ran}(R)^{\perp},~\langle\eta,e\rangle\neq 0]=0$, which shows that it holds with full probability that $\eta\in(\mathrm{ran}(R)^{\perp})^{\perp}=\mathrm{cl}(\mathrm{ran}(R))$, concluding the proof. ∎
Proof of Corollary E.20.
Under the current assumptions, the scaled noise $\bar{\eta}_{n}=(R^{\sharp})^{1/2}\eta_{n}$ is $1$-subgaussian conditionally on $\mathcal{F}$, where $R^{\sharp}$ is the Moore-Penrose pseudo-inverse of $R$. Indeed, it is (conditionally) 0-mean and for any $g\in\mathcal{G}$,
@@
The second equality comes from the fact that $R^{1/2}$ and $R^{\sharp}$ commute; the proof is a technical exercise that follows from the spectral theorem applied to $R_{\mathcal{V}}$ and the spectral definition of $R^{\sharp}$. We also leveraged in the last inequality the fact that $RR^{\sharp}$ is an orthogonal projection, and thus $\lVert g\rVert_{RR^{{\sharp}}}\leq\lVert g\rVert_{\mathcal{G}}$; see for instance Proposition 2.3 in Engl et al. [45]. Theorem E.14 applied with $Q=R_{\mathcal{V}}$ then guarantees that for all $\delta\in(0,1)$, it holds with probability at least $1-\delta$ that, for all $N\in{\mathbb{N}_{>0}}$,
$\lVert\iota_{\mathcal{H}}^{-1}M_{N}^{\star}(R^{\sharp}_{N})^{1/2}\eta_{:N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes R_{\mathcal{V}}}\leq\sqrt{2\mathrm{Tr}(R_{\mathcal{V}})\ln\left(\frac{1}{\delta}\left[\det(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}^{\star}\tilde{V}^{-1}\tilde{M}_{N})\right]^{1/2}\right)}.$
@@
$\lVert\iota_{\mathcal{H}}^{-1}M_{N}^{\star}(R^{\sharp}_{N})^{1/2}\eta_{:N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes R_{\mathcal{V}}}=\rho^{-1}\lVert M_{N}^{\star}\eta_{:N}\rVert_{(V+V_{N})^{-1}},~\text{\acs{as}}$
@@
$\qquad\qquad\left.(\tilde{L}_{m}^{\star}\otimes\mathrm{id}_{\mathcal{V}})\iota_{\mathcal{G}}^{-1}\eta_{m}\right\rangle_{{\tilde{\mathcal{H}}}\otimes\mathcal{V}}$
@@
where we leveraged in the third equality the fact that $(R_{\mathcal{V}}^{\sharp})^{1/2}$ and $R_{\mathcal{V}}$ commute, and that $(\tilde{V}+\tilde{V}_{N})^{-1}\otimes\mathrm{id}_{\mathcal{V}}=\iota_{\mathcal{H}}^{-1}(V+V_{N})^{-1}\iota_{\mathcal{H}}$ in the fourth one. Next, $RR^{\sharp}$ is the orthogonal projection on $\mathrm{cl}(\mathrm{ran}(R))$; see for instance Proposition 2.3 in [45]. Yet, $\eta_{n}$ is a.s. in $\mathrm{cl}(\mathrm{ran}(R))$ by Lemma E.22. It follows that $RR^{\sharp}\eta_{n}=\eta_{n}$, a.s., for all $n\in{\mathbb{N}_{>0}}$. As a result, it holds a.s. that
@@
$=\rho^{-2}\lVert M_{N}^{\star}\eta_{:N}\rVert_{(V+V_{N})^{-1}}^{2}.$